EmilyHerron:

A machine cuts a strip of carpet into two pieces. The length of the smaller piece is 5 meters greater than the length of the larger piece. If the length of the smaller piece is 12 meters, the length of the bigger piece is blank meters and the total length of the carpet is blank meters.

AZ:

By the way, closing the question and re-posting it doesn't like increase the chances of it being answered haha The bump button does because it can push it to the top but your post has already been at the top :b

EmilyHerron:

wow how stupid am i lol🤦

AZ:

So for this question, do you know we can write the equations? Length of smaller piece is 5 meters greater than the length of the larger piece Length of smaller piece is 12 what is the length of the larger piece? and what is the length of the total length of the carpet?

AZ:

No no no, you're new so I wouldn't expect you to know. Welcome to QuestionCove by the way!

EmilyHerron:

Thanks

AZ:

So do you know how to write it as an equation so that way we can solve it?

EmilyHerron:

no stuck math is not my thing

AZ:

So we have Length of smaller piece is 5 meters greater than the length of the larger piece Let's use some letters like S for small piece and L for large piece Following me still?

EmilyHerron:

yep

AZ:

So we have S is 5 meters greater than L now does the word greater than sound like addition, multiplication, division, or subtraction?

EmilyHerron:

multiplication:/

AZ:

nooo

EmilyHerron:

Subtraction

EmilyHerron:

1 attachment
AZ:

@EmilyHerron no

AZ:

It's addition

1 attachment
AZ:

think of the word 'is' as an equal sign so we have S is 5 meters greater than L S = 5 greater than L change `greater than` to a plus sign what do you get

EmilyHerron:

S=5+L

AZ:

Great! Now remember they told you that the length of the smaller piece (S) is 12 so S = 5 + L 12 = 5 + L what is L?